# Simulation of the continuous-time random walk using subordination schemes

### v1

@@ II.1 Model
with $\xi=(N-t/\langle\tau\rangle)/(t/\bar{t})^{1/\alpha}$ and $\bar{t}=\langle\tau\rangle^{1+\alpha}/(\tau_{0}^{\alpha}|\Gamma(1-\alpha)|)$. Eq. (5) is plotted by the red solid line in Fig. 1. Here the symbol $\mathcal{L}_{\alpha}(\xi)$ is the Lévy stable distribution, whose Fourier transform is $\exp((ik)^{\alpha})$. Typical fluctuations Eq. (5) describes the case when $N-t/\langle\tau\rangle$ is of the order of $t^{1/\alpha}$, i.e., the central part of the distribution. Using the asymptotic behavior of the Lévy distribution, the far left tail of $N$ follows $Q_{t}(\xi)\sim(-\xi)^{-1-\alpha}/\Gamma(-\alpha)$. It indicates that Eq. (5) does not give the information of the MSD Wang2018Renewal .
@@ III.1 Typical fluctuations
$P(\xi)\sim\mathcal{L}_{\alpha}(\xi)$ (17)
with $\xi=(x-at/\langle\tau\rangle)/l(t)\rangle$ and $l(t)=a(t/\overline{t})^{1/\alpha}$. In Eq. (17), it is crucial to emphasize that $a$ must be nonzero; otherwise, the positional distribution would yield a Gaussian distribution. In the general scenario, $P(x,t)$ is described by the convolution of a Lévy stable distribution and a Gaussian distribution Wang2020Fractional
@@ III.2 Rare fluctuations
Figure 8: Mean of the position under various biases are depicted, with $\Box$’ and ’$\times$’ representing simulations and theoretical predictions, respectively. Solid lines correspond to the asymptotic prediction given by Eq. (21). The parameters are $\sigma=1$, $\alpha=3/2$, $\tau_{0}=0.1$, and $2\times 10^{7}$ realizations.
@@ III.3 Moments of the position
$\langle\hat{x}^{q}(s)\rangle=(-i)^{q}\frac{\partial^{q}\widetilde{\widehat{P}}(k,s)}{\partial k^{q}}|_{k=0},$
@@ III.4 Breakthrough curves
$P(x,t)\sim\int_{-\infty}^{\infty}\frac{1}{\sqrt{4\pi c_{m1}}}\exp\left[-\frac{(x-y-c_{m2})^{2}}{4c_{m1}}\right]\frac{1}{(c_{m3})^{1/\beta}}\mathcal{L}_{\beta}\left[\frac{y}{(c_{m3})^{1/\beta}}\right]{\rm d}y$ (23)
@@ Acknowledgments
WW expresses gratitude to Eli Barkai for the discussion that inspired this work. The work is supported by the National Natural Science Foundation of China under Grant No. 12105243 and the Zhejiang Province Natural Science Foundation LQ22222222A050002050002050002050002.

### v2

@@ II.1 Model
[…]
@@ III.1 Typical fluctuations
$P(\zeta)\sim\mathcal{L}_{\alpha}(\zeta)$ (17)
with $\zeta=(x-at/\langle\tau\rangle)/l(t)\rangle$ and $l(t)=a(t/\overline{t})^{1/\alpha}$. In Eq. (17), it is crucial to emphasize that $a$ must be nonzero; otherwise, the positional distribution would yield a Gaussian distribution. In the general scenario, $P(x,t)$ is described by the convolution of a Lévy stable distribution and a Gaussian distribution Wang2020Fractional
@@ III.2 Rare fluctuations
Figure 8: Mean of the position under various biases are depicted. Lines correspond to the asymptotic prediction given by Eq. (21). The parameters are $\sigma=1$, $\alpha=3/2$, $\tau_{0}=0.1$, and $2\times 10^{7}$ realizations.
@@ III.3 Moments of the position
$\langle\hat{x}^{q}(s)\rangle=i^{q}\frac{\partial^{q}\widetilde{\widehat{P}}(k,s)}{\partial k^{q}}|_{k=0},$
@@ III.4 Breakthrough curves
$P(x,t)\sim\int_{-\infty}^{\infty}\frac{1}{\sqrt{4\pi c_{m1}}}\exp\left[-\frac{(x-y-c_{m2})^{2}}{4c_{m1}}\right]\frac{1}{(c_{m3})^{1/\alpha}}\mathcal{L}_{\alpha}\left[\frac{y}{(c_{m3})^{1/\alpha}}\right]{\rm d}y$ (23)
@@ Acknowledgments
WW expresses gratitude to Eli Barkai for the discussion that inspired this work. The work is supported by the National Natural Science Foundation of China under Grant No. 12105243 and 12101555, and the Zhejiang Province Natural Science Foundation LQ22222222A050002050002050002050002.